# Safe MPC Alignment with Human Directional Feedback

### v2

@@ VI-A2 Convergence Analysis
With each simulated human correction, we track $\boldsymbol{\theta}_{i}$ in each learning update. Each update of ${\Theta}_{i}$ takes less than 30ms. $\boldsymbol{\theta}_{i}$ falls in the ground-truth set $\bar{\Theta}_{H}$ after 14 corrections. The MPC rollout trajectory (phase plot) with each $\boldsymbol{\theta}_{i}$ is shown in Fig. 8. The ground-truth constraint (27) and the learned constraints $\boldsymbol{g}_{\boldsymbol{\theta}_{i}}$ are also plotted in Fig. 8. From Fig. 8, we can conclude that the ground-truth constraint can be successfully learned from simulated corrections.
(a) Initial trajectory (with $\boldsymbol{\theta}_{1}$)
(b) After 3 corrections (with $\boldsymbol{\theta}_{4}$)
(c) After 6 corrections (with $\boldsymbol{\theta}_{7}$)
(d) After 14 corrections (with $\boldsymbol{\theta}_{15}$)
Figure 8: The pendulum rollout trajectory at different correction interactions: (a) trajectory with $\boldsymbol{\theta}_{1}$, (b) with $\boldsymbol{\theta}_{4}$, (c) with $\boldsymbol{\theta}_{7}$, (d) with $\boldsymbol{\theta}_{15}$. The learned constraint with $\boldsymbol{\theta}_{1}$ is omitted since it covers the whole first quadrant. As the learning proceeds, the learned constraint boundary (orange line) converges to the ground-truth constraint boundary (red line). When the learning process terminates, the learned and ground-truth constraint boundaries almost overlap, suggesting the success of constraint learning from corrections.
To further analyze the learning convergence, we run the algorithm for 10 times. In each run, we keep track of two metrics over correction (iteration) count and show them in Fig. 9. The first metric (left panel) is the logarithm volume measure of the solved MVE $\log\det\bar{H}$ in (20); this provides a lower bound of $\Theta_{i}$ to show the exponential convergence property. The second (right panel) is the parameter error $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$. From Fig. 9, we can conclude the proposed method is
  • (I)
    Fast-converging. The logarithm volume $\log\det\bar{H}$ has near-linear decreasing speed. This is consistent with the exponential convergence in theory stated in Lemma 5.
  • (II)
    Correction-efficient. $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$ converges to the value lower than 0.02 within 17 corrections. This is consistent with the theoretical upper bound of correction count from given by Theorem 1 :
@@
Figure 9: (a) The plot of $\log\det\bar{H}$ with respect to correction numbers. In all algorithm runs, the plot shows the near-linear decrease of the logarithm volume, suggesting the exponential decrease of the volume of MVE of the hypothesis spaces, as asserted in Theorem 1. (b) The plot of the parameter estimation error $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$ versus correction numbers. In all runs, the parameter estimation error can converge to a value lower than 0.02 within 17 corrections.
VI-A3 Analysis for Misspecification Case
To analyze the performance of our algorithm to certify misspecification cases, we set the initial hypothesis space $\Theta_{0}$ as ${\boldsymbol{c}_{l}}=[-1,-1]^{\mkern-1.5mu\mathsf{T}}$ and ${\boldsymbol{c}_{u}}=[0.8,0.8]^{\mkern-1.5mu\mathsf{T}}$ which does not contain $\bar{\Theta}_{H}$. In this setting, the algorithm automatically stops after 7th correction. The path of $\boldsymbol{\theta}_{i},\ i=1,\dots,8$ is shown in Fig. 10. In Fig. 10, $\boldsymbol{\theta}_{i}$ is continuously moving towards the boundary of $\Theta_{0}$. This fact verifies Theorem 2 and justifies the proposed method is capable of misspecification detection.
Figure 10: Illustration of misspecification case. The blue dashed lines are the boundary of $\Theta_{0}$. $\boldsymbol{\theta}_{i},\ i=1,\dots,7$ is shown in he orange dots in one run. Black arrows show the update directions of $\boldsymbol{\theta}_{i}$s. The center of the ground-truth intent set is marked as the red star, and the $\boldsymbol{\theta}_{8}$ at the termination of learning is marked as the green triangle. When misspecification happens, $\boldsymbol{\theta}$ will converge toward the boundary of $\Theta_{0}$, as stated by Theorem 2.
VI-B Two-link Robot Arm Reaching with Complex Constraint
@@ VI-B1 Setup
The dynamics if torque controlled two-link robot arm moving in the horizontal plane is $M(\boldsymbol{q})\ddot{\boldsymbol{q}}+\boldsymbol{c}(\boldsymbol{q},\dot{\boldsymbol{q}})=\boldsymbol{\tau}$, where $M(\boldsymbol{q})$ is the inertia matrix; $\boldsymbol{c}(\boldsymbol{q},\dot{\boldsymbol{q}})$ is the Coriolis term. $\boldsymbol{q}=[q_{1},q_{2}]^{\mkern-1.5mu\mathsf{T}}$ defines the two joint angles of the robot. The system state is defined as $\boldsymbol{x}=[\boldsymbol{q},\dot{\boldsymbol{q}}]$. The control input $\boldsymbol{u}$ is the joint torque $\boldsymbol{\tau}=[\tau_{1},\tau_{2}]^{\mkern-1.5mu\mathsf{T}}$, i.e. $\boldsymbol{u}=\boldsymbol{\tau}$. The other detailed physical parameters are given in Appendix.
As shown in Fig. 12, the robot starts at $\boldsymbol{x}_{\mathrm{init}}=[-\frac{\pi}{2},0,0,0]^{\mkern-1.5mu\mathsf{T}}$ and the task is to reach a target $\boldsymbol{x}^{*}=[\frac{\pi}{2},0,0,0]^{\mkern-1.5mu\mathsf{T}}$. In Safe MPC (1), we set $T=30$, the step cost functions
$\begin{gathered}c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})=\boldsymbol{x}_{t}^{\mkern-1.5mu\mathsf{T}}P\boldsymbol{x}_{t}+\boldsymbol{u}_{t}^{\mkern-1.5mu\mathsf{T}}R\boldsymbol{u}_{t},\\ h(\boldsymbol{x}_{T})=(\boldsymbol{x}_{T}-\boldsymbol{x}^{*})^{T}Q(\boldsymbol{x}_{T}-\boldsymbol{x}^{*}),\end{gathered}$ (31)
with $P=\mathrm{diag}(1,1,2,2)$, $R=\mathrm{diag}(0.1,0.1)$ and $Q=\mathrm{diag}(5,5,10,10)$.
As shown by the black lines (obstacles) in Fig. 12, we define a complex reaching environment with non-convex obstacles and narrow corridor. The ground-truth global safety constraint corresponding to Fig. 12 is given
${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi})=-1+\boldsymbol{\theta}_{H}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ (32)
with $\boldsymbol{\phi}{=}[\mathrm{tanh}(x_{e}y_{e}),\mathrm{tanh}(x_{e}^{2},\mathrm{tanh}(y_{e}^{2}),\mathrm{tanh}(x_{e}),\mathrm{tanh}(y_{e})]^{\mkern-1.5mu\mathsf{T}}$, $\boldsymbol{\theta}_{H}=[0.78,0.91,-0.52,0.52,-1.04]^{\mkern-1.5mu\mathsf{T}}$, and $[x_{e},y_{e}]^{\mkern-1.5mu\mathsf{T}}$ is the end-effector position of the robot arm. The human intent set is $\bar{\Theta}_{H}{=}\{\boldsymbol{\theta}|\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.2\}$. We set the initial hypothesis $\Theta_{0}$ with box boundaries $\boldsymbol{c}_{l}=[-3,\dots,-3]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[3,\dots,3]^{\mkern-1.5mu\mathsf{T}}$. The simulated human corrections are generated in the similar way to that in Section VI-A.
Figure 11: The heatmap of a learned constraint $g$. The learned and ground-truth constraints are almost identical, showing the success of our method for learning complex constraint from corrections.
@@ VI-B2 Result
Our method can successfully learn the safety constraint within 25-30 simulated corrections. One instance of the learned constraint is shown in Fig. 11. The comparison of the robot reaching motion without and with the learned safety constraint is shown in Fig. 12. The robot is capable of operating in the intended safe region after the algorithm converges. Those results show that the efficacy of the proposed method for solving system with complex safety constraints.
(a) Trajectory Before Corrections
(b) Trajectory After Corrections
Figure 12: The trajectory of the two-link robot arm (a) without and (b) with learned safety constraint. The robot successfully achieves the task with the learned safety constraint.
@@ VI-C Baseline Comparison
In this session, we compare our method with two baseline methods using magnitude corrections in the pendulum environment (the task and model settings are identical to Section VI-A). As reviewed previously, the direct literature learning safety constraints from human feedback is sparse to the authors’ best knowledge. Thus, the chosen baselines are an adaptation from the most relevant methods.
Baseline 1: gradient matching. This baseline is inspired by [59], where raw human correction (with magnitude) is directly used as the gradient direction. In a similar spirit, when a human feedback $\mathbf{a}$ is made on robot trajectory $\xi_{\boldsymbol{\theta}}^{\gamma}$, we directly use the human magnitude correction as the true gradient label to learn the constraint $\boldsymbol{\theta}$. This leads to a learning loss
$\mathcal{L}=\left\lVert\mathbf{a}+\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\right\rVert^{2}$ (33)
Baseline 2: maximum likelihood learning. This baseline follows [60], where it views $e^{{g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})}$ as a probability of human applying the "stop-feedback" to robot motion $\boldsymbol{\xi}$, when the robot reaches in a $\epsilon_{g}$-proximity of the safety boundary, i.e., when ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ Thus, a dataset of trajectory-correction paris, i.e., $\mathcal{D}=\{\mathbf{a}_{i},\boldsymbol{\xi}_{i}\}_{i=0}^{N}$, can be used as the training data to learn $\boldsymbol{\theta}$ using maximum likelihood, while also penalizing positivity of ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i})$ as a regulizer term. This leads to the following loss
$\mathcal{L}=\sum\nolimits_{i=1}^{N}\big{(}-{g}_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i})+\alpha\cdot\mathrm{softplus}(g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i}))\big{)}$ (34)
In our comparison, we choose $N=15$ (similar to correction number needed for our method to converge) and $\alpha=5$ for its best performance.
The comparison between our method and Baseline 1 focuses on data efficiency. The convergence criteria of $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.02$ is set for both methods. 50 independent runs are performed and the number of correction needed for convergence are counted. We use Adam optimizer [61] with learning rate 0.02. We summarize the results in Table I. The parameter estimation error $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ versus learning iteration (correction count) for a best run (i.e., best performance for the baseline) is plotted in Fig. 13(a). The comparison results demonstrate the significant data efficiency of the proposed method compared to gradient matching baseline, even though the latter uses both magnitude and direction information of corrections to learn.
@@
(a) Comparison with Baseline 1
(b) Comparison with Baseline 2
Figure 13: (a) Comparison with gradient matching baseline: the parameter estimation error $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ versus learning iteration (correction count). This is best run (among 50 runs) with the least corrections for the baseline. Our method converges after 12 corrections, while the baseline needs more than 160 corrections to converge. (b) Comparison with the maximum likelihood baseline under the different $\epsilon_{g}$. In the plot, the y-axis is $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ at convergence and the error bar denotes the standard deviation. The maximum likelihood baseline fails to learn the correct weights when $\epsilon_{g}$ is large, while our method consistently shows high learning accuracy.
The comparison between our method and maximum likelihood baseline focuses on the varying positions of simulated correction, i.e., varying $\epsilon_{g}$. Recall that the simulated corrections will be generated only when ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$. The value of $\epsilon_{g}$ determines the position of the correction applied. In our comparison, we vary $\epsilon_{g}$ from 0.1 to 0.9 and track the converged parameter error for both methods. As shown in Fig. 13(b), safe MPC alignment consistently converges to the ground-truth $\boldsymbol{\theta}_{H}$ regardless of $\epsilon_{g}$, while the maximum likelihood baseline fails when $|\epsilon_{g}|$ is large. The results suggest the robustness of the proposed method against different correcting conditions compared to maximum likelihood baseline method.

### v3

@@ VI-A2 Convergence Analysis
Figure 8: (a) The plot of $\log\det\bar{H}$ with respect to correction numbers. In all algorithm runs (different colors), the plot shows the near-linear decrease of the logarithm volume, suggesting the exponential decrease of the volume of MVE of the hypothesis spaces, as asserted in Theorem 1. (b) The plot of the parameter estimation error $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$ versus correction numbers. In all runs, the parameter estimation error can converge to a value lower than 0.02 within 17 corrections, consistent with the theoretical bound in Theorem 1.
With each simulated correction, we track $\boldsymbol{\theta}_{i}$ in each learning update. Each update of ${\Theta}_{i}$ takes less than 30ms. $\boldsymbol{\theta}_{i}$ falls in the ground-truth set $\bar{\Theta}_{H}$ after 14 corrections. The MPC rollout trajectory (phase plot) with each $\boldsymbol{\theta}_{i}$ is shown in Fig. 7. The ground-truth constraint (27) and the learned constraints $\boldsymbol{g}_{\boldsymbol{\theta}_{i}}$ are also plotted. Fig. 7 shows that the true safety constraint can be successfully learned from simulated correction.
To analyze the convergence, we run the algorithm for 10 times. In each run, we keep track of two metrics over corrections (iterations) and show them in Fig. 8. The first metric (left panel) is the logarithm volume measure of the solved MVE $\log\det\bar{H}$ in (20); this provides a lower bound of the volume of $\Theta_{i}$ to show the exponential convergence. The second (right panel) is the parameter error $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$. From Fig. 8, we can conclude the proposed method is
  • (I)
    Fast-converging. The logarithm volume $\log\det\bar{H}$ has a near-linear decreasing speed. This is consistent with the theory in Lemma 5.
  • (II)
    Data-efficient. $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$ converges to the value lower than 0.02 within 17 corrections. This is consistent with the theoretical upper bound of correction count in Theorem 1:
VI-A3 Misspecification Analysis
Figure 9: Experiment of misspecification case. The blue dashed lines are the boundary of $\Theta_{0}$. Path of $\theta$ convergence, $\boldsymbol{\theta}_{i},\ i=1,\dots,7$, is shown in orange dots with arrows showing the update direction. The ground-truth intent set is marked as the red star. Since misspecification happens for this case, $\boldsymbol{\theta}$ will converge toward the boundary of $\Theta_{0}$ (the green triangle). This is consistent with the statement in Theorem 2.
To analyze the performance of our algorithm to certify misspecification cases, we set the initial hypothesis space $\Theta_{0}$ as ${\boldsymbol{c}_{l}}=[-1,-1]^{\mkern-1.5mu\mathsf{T}}$ and ${\boldsymbol{c}_{u}}=[0.8,0.8]^{\mkern-1.5mu\mathsf{T}}$ which does not contain $\bar{\Theta}_{H}$. In this setting, the algorithm automatically stops after 7th correction. The path of $\boldsymbol{\theta}_{i},\ i=1,\dots,8$ is shown in Fig. 9. $\boldsymbol{\theta}_{i}$ is continuously moving towards the boundary of $\Theta_{0}$. This fact verifies Theorem 2 and justifies the proposed method is capable of misspecification detection.
VI-B Quadrotor Navigation in Complex Environment
Figure 10: Quadrotor navigation in an 3D zigzag tube.
@@ VI-B1 Setup
We consider a thrust-controlled quadrotor navigating in an 3D zigzag tube, shown in Fig. 10. The simulation environment is based on MuJoCo [todorov2012mujoco], and the quadrotor dynamics follows [jin2022learning]. Initially, the quadrotor has no knowledge of the safety constraint. When it nears the tube wall, a simulated (synthetic) correction is generated to push it away from the tube wall. Details of the setting are provided in Appendix B-B.
To learn positional safety constraints, we experiment with three feature vectors $\boldsymbol{\phi}(\boldsymbol{\xi})$ in (5):
Polynomial feature vector
A vector of polynomial basis functions up to 3rd order to properly capture the zig-zag pattern of the tube) is defined in the quadrotors’ 3D position $[x,y,z]$ at the 5th timestep in motion plan $\boldsymbol{\xi}$:
$\boldsymbol{\phi}(\boldsymbol{\xi}){=}[x_{5}^{3},y_{5}^{3},z_{5}^{3},x_{5}^{2},y_{5}^{2},\\ z_{5}^{2},x_{5},y_{5},z_{5},x_{5}y_{5},y_{5}z_{5},x_{5}z_{5}]^{\top}.$ (31)
The features are defined using the quadrotor’s 3D position at the 5th prediction step. The hypothesis space is initialized as a box space with $\boldsymbol{c}_{l}=[-80,\dots,-80]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[200,\dots,200]^{\mkern-1.5mu\mathsf{T}}$. We select $\phi_{0}=1$ and $\gamma=60$.
Pre-trained neural feature vector
We obtain the neural spacial features from a pre-trained Signed Distance Function (SDF) [park2019deepsdf]. Specifically, we pre-train an SDF neural network using the point cloud data samples near the tube surface. This simulates the application where some geometric information of the environment is available. The neural SDF is pre-trained in a regression-like manner, with distance, normal, and Eikonal losses (see details in Appendix B-B). With the pretrained SDF function, we take the 16-dimensional output of the second-to-last layer of the SDF network as our feature vector $\boldsymbol{\phi}(\boldsymbol{\xi})$. The input to $\boldsymbol{\phi}(\boldsymbol{\xi})$ is the quadrotor’s 3D position $[x,y,z]$ at the 5th perdition step. Such a choice of feature vector $\boldsymbol{\phi}(\boldsymbol{\xi})$ means that we will use the proposed Safe MPC Alignment to fine tune the last layer of the learned SDF, $g(\boldsymbol{\xi})=\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi})$, for its safe navigation control.
To facilitate the real-time Safe MPC with the neural safety constraint $g(\boldsymbol{\xi})=\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi})$, the optimization (1) uses the linearized $g(\boldsymbol{\xi})$ for calculation, which is linearized at the real robot state $\boldsymbol{x}^{\text{real}}_{k}$. The initial bounds of the hypothesis space are set to $\boldsymbol{c}_{l}=[-0.2,\dots,-0.2]^{\mkern-1.5mu\mathsf{T}}$, $\boldsymbol{c}_{u}=[0.1,\dots,0.1]^{\mkern-1.5mu\mathsf{T}}$. We select $\phi_{0}=0.1$ and $\gamma=50$.
To ensure the feasibility of the learned constraint, in the above feature vectors, two additional cuts are added in each MVE update to enforce that the start and goal remain within the intended safe region. The task is successful if the drone reaches the goal without triggering this stop.
Polynomial features with velocity features
We consider the safety constraints are additionally defined in the velocity space. After learning the positional constraint with the polynomial feature vector, a new set of velocity-based features over all prediction steps, $\boldsymbol{\phi}(\boldsymbol{\xi})=[\|\boldsymbol{v}_{1}\|^{2},\dots,\|\boldsymbol{v}_{10}\|^{2}]^{\top}$, is added. In this case, we will freeze the weights for polynomial features, and only learn the weights for velocity features. Our objective is to impose an additional constraint that caps the quadrotor’s velocity magnitude below 0.45 [m/s], while ensuring that the previously learned positional constraint remains satisfied for collision avoidance. When the quadrotor’s velocity exceeds a correction triggering threshold (0.4), a simulated correction is applied to push the quadrotor in the direction of negative velocity. The hypothesis space for the velocity features is initialized with bounds $\boldsymbol{c}_{l}=[0,\dots,0]^{\top}$ and $\boldsymbol{c}_{u}=[0.16,\dots,0.16]^{\top}$. We set the bias feature $\phi_{0}=1$ and select $\gamma=100$.
Figure 11: The Safe MPC rollout trajectories (in green) for learning safety constraints with polynomial features after 26, 38, 45 corrections (different columns). The first row and second row show the top and the side views of the navigation environment.
Figure 12: The Safe MPC rollout trajectories (in green) for learning safety constraints with neural features after 6, 9, 11 corrections (different columns). The first row show the top view, and second row shows the side view of the navigation environment.
@@ VI-B2 Result
Fig. 11 and Fig. 12 show the rollout trajectories of the learned Safe MPC policy using polynomial and neural features, respectively. With polynomial features, the task is completed in 45 corrections, while the neural features require only 11. This is because the neural features, constructed from the second-to-last layer of pre-trained neural SDF, have already encoded the partial knowledge of spatial constraints.
Figure 13: Velocity profile during the first 100 timesteps (acceleration phase) of the drone trajectory. After 12 corrections, the quadrotor successfully learns to stay below the intended velocity cap of 0.45 (dashed line).
For the learning of velocity constraint, Fig. 13 shows the velocity profile of the quadrotor during the Safe MPC rollout in the 100 time steps (acceleration phase) after different number of simulated corrections (different colors). After 12 corrections, the quadrotor successfully learns to stay below the intended velocity cap of 0.45 (dashed line). Throughout this process, the positional safety constraint remains satisfied, with the quadrotor continuing to navigate within the tube to the target. This result demonstrates the flexibility of the proposed method in scenarios involving both positional and velocity constraints.
@@ VI-C Baseline Comparison
In this session, we compare our method with two baseline methods using magnitude corrections in the pendulum environment. Since the literature on directly learning safety constraints from human feedback is sparse, the chosen baselines are an adaptation from the most relevant methods.
VI-C1 Baseline 1: Gradient Matching
This baseline is inspired by [losey2019learning], where raw human correction (with magnitude) is directly used as the gradient direction. In a similar spirit, when a human correction $\mathbf{a}$ is made on robot motion $\xi_{\boldsymbol{\theta}}^{\gamma}$, we directly use the human correction as the true gradient label to learn the constraint $\boldsymbol{\theta}$. This leads to a learning loss
$\mathcal{L}(\boldsymbol{\theta})=\left\lVert\mathbf{a}+\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta})\right\rVert^{2}.$ (32)
VI-C2 Baseline 2: Maximum Likelihood
. This baseline follows [poletti2023learning], where it views $e^{{g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})}$ as a probability of human applying the "stop-feedback" to robot motion $\boldsymbol{\xi}$, when the robot reaches in a $\epsilon_{g}$-proximity of the safety boundary, i.e., when ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ . Thus, a dataset of trajectory-correction paris, i.e., $\mathcal{D}=\{\mathbf{a}_{i},\boldsymbol{\xi}_{i}\}_{i=0}^{N}$, can be used as the training data to learn $\boldsymbol{\theta}$ using maximum likelihood, while also penalizing positivity of ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i})$ as a regulizer term. This leads to the following loss
$\mathcal{L}(\boldsymbol{\theta})=\sum\nolimits_{i=1}^{N}\big(-{g}_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i})+\alpha\cdot\mathrm{softplus}(g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i}))\big).$ (33)
We choose $N=15$ (similar to the correction number needed for our method to converge) and $\alpha=5$ for its best performance.
The comparison between our method and the gradient matching baseline focuses on data efficiency. The convergence criteria of $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.02$ is set for both methods. 50 independent runs are performed and the total number of correction at convergence are counted. We use Adam optimizer [diederik2014adam] with a learning rate of 0.02. We summarize the results in Table I. The parameter error $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ versus learning iteration for a best run (for the baseline) is plotted in Fig. 14(a). The comparison results demonstrate data efficiency of the proposed method compared to the gradient matching baseline, even though the latter uses both magnitude and direction information of corrections to learn.
@@ VI-C2 Baseline 2: Maximum Likelihood
(a) Comparison: gradient matching
(b) Comparison: maximum likelihood
Figure 14: (a) Comparison with gradient matching baseline: the parameter error $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ versus learning iteration. This is best run with the fewest corrections for the baseline. Our method converges after 12 corrections, while the baseline needs more than 160 corrections. (b) Comparison with the maximum likelihood baseline under the different $\epsilon_{g}$: the y-axis is $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ at convergence with the error bar denoting the standard deviation. The maximum likelihood baseline fails to learn the correct weights when $\epsilon_{g}$ is large.
The comparison with maximum likelihood baseline uses the varying positions of simulated correction, i.e., varying $\epsilon_{g}$. Recall that the simulated corrections are generated only when ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$. The value of $\epsilon_{g}$ determines the position of correction. In our comparison, we vary $\epsilon_{g}$ from 0.1 to 0.9 and track the converged parameter error for both methods. As shown in Fig. 14(b), Safe MPC alignment consistently converges to the true $\boldsymbol{\theta}_{H}$ regardless of $\epsilon_{g}$, while the maximum likelihood baseline fails when $|\epsilon_{g}|$ is large. The results suggest the robustness of the proposed method against different correcting conditions compared to the maximum likelihood baseline.